# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v2

@@ 8.1 Examples and video source details
  • SoccerNet [giancola2018soccernet] – 56 videos were selected from this dataset, each averaging 92.4 minutes in duration. These structured soccer match recordings include 389 questions. The mean GT time interval is 61.9 seconds.
  • MovieChat-1K Films [song2024moviechat] – a total of 140 film clips, each 8 minutes long, were selected from the dataset. Clips from the same film were combined to create 24 film segments, with an average duration of 46.4 minutes each. This subset contains 122 questions, with a mean GT time interval of 22.4 seconds.
  • Walking Tours Dataset [venkataramanan2023imagenet] – 12 high-resolution videos ($3840\times 2160$) were selected from this dataset, averaging 81.3 minutes. These videos capture city tours from an egocentric perspective and include 65 questions. The mean GT time interval is 31.4 seconds.
@@ 8.3 Localization evaluation
FALCON-Bench requires models to provide an evidence of the answer (short clip) rather than precisely matching the entire clip where the answer is located. To achieve this goal, we leverage The Ground Truth over Union (GToU) metric to compare the predicted and the GT temporal interval. Unlike the commonly used Intersection over Union (IoU) in temporal grounding tasks [ding2023temporal], GToU assigns a score of 1.0 if the predicted interval is entirely within the GT interval, regardless of the degree of overlap. In all other cases, GToU behaves identically to IoU (Figure 5). In mathematical terms,
@@ 10.1 FALCON-Bench
mPLUG-Owl3 [ye2024mplug] 57.6 39.2 21.1 26.0 28.7 2.24 1.20 1.58 1.67
LLaVA-OV [llava-onevision] 72.2 45.0 28.5 46.0 39.8 2.46 1.57 2.42 2.15
@@
VideoChat2-HD [li2024mvbench] 27.2 18.6 12.1 22.0 17.5 1.24 0.67 1.34 1.08
Video-LLAVA [lin2023video] 34.0 19.6 6.49 18.0 14.6 1.20 0.46 1.14 0.93
LLaVA-Video [llava-next] 79.2 61.7 44.0 52.0 52.6 3.31 2.38 2.74 2.81
@@
MovieChat-OV [song2024moviechat] 44.5 5.88 18.3 14.0 12.7 0.49 1.02 0.82 0.77
Apollo [zohar2024apollo] 71.3 50.9 31.3 48.0 43.4 2.77 1.79 2.62 2.39
@@
mPLUG-Owl3 [ye2024mplug] 82.9 59.8 53.1 58.0 65.2 3.25 2.89 3.18 3.10
LLaVA-OV [llava-onevision] 84.4 40.1 10.4 54.0 34.8 2.02 0.53 2.80 1.78
@@ 10.3 VLMs Calibration
To measure VLM calibration, we adopt the Reliability Diagrams [degroot1983comparison]. These diagrams group all the predictions in bins according to their confidence and measure the gap between confidence and accuracy per each bin. Specifically, we split the $N$ predictions in $M=10$ bins according to their confidence and, for each bin $B_{m}$, we compute its count $N_{m}$ average confidence $C_{m}$ and its average accuracy $A_{m}$. From these diagrams, we may compute the Average Calibration Error (ACE) as,
@@
where $M^{+}$ is the number of non-empty bins [neumann2018relaxed]. We compute Calibration Count (CC), which quantifies the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE). For example, the Calibration Count at threshold 0.9 is computed as,
@@
We evaluated various confidence aggregation methods, specifically likelihood, average, and geometric average, as done in similar calibration studies with LLMs [liu2023litcab]. Looking first at the reliability diagrams (Figure 10 (a-f)), we discard likelihood as a suitable confidence metric. The distribution of confidence values is highly spread out, with many answers assigned extremely low confidence. This occurs because likelihood multiplies the probability of all tokens, thus longer answers tend to receive lower confidence scores, making it unreliable for calibration.
@@ 10.5 FALCON-Bench vs QAEgo4D
The QAEgo4D [Baermann_2022_qaego4d] dataset is the closest to FALCON-Bench in terms of the task definition as it also addresses the problem of VAS, that the authors call question answering visual language grounding (VLG). They also have open ended questions. However, besides the time differences in the video durations –FALCON-Bench with an average duration of $\sim 80$ minutes while QAEgo4D has an average duration of $\sim 8$ minutes–, the type of questions are completely different. QAEgo4D questions are automatically generated from the sparse video narration, focusing on the main object or action: Q: What did I Put in the Pan?A: cheese, or Q: What paint can did I open?A:black paint. Meanwhile, FALCON-Bench questions are human curated to be challenging: Q: What is the number of the train that crosses paths with Lightning McQueen at night?A:A113; Q:What message about the flu appears on a city building?A:Bovril nourishes you to resist ’flu or Q:Which player of Chelsea received a red card?A: Thibaut Courtois.

### v3

@@ 9.1 Examples and video source details
  • SoccerNet [20] – 56 videos were selected from this dataset, each averaging 92.4 minutes in duration. These structured soccer match recordings include 389 questions. The mean GT time interval is 61.9 seconds.
  • MovieChat-1K Films [45] – a total of 140 film clips, each 8 minutes long, were selected from the dataset. Clips from the same film were combined to create 24 film segments, with an average duration of 46.4 minutes each. This subset contains 122 questions, with a mean GT time interval of 22.4 seconds.
  • Walking Tours Dataset [48] – 12 high-resolution videos ($3840\times 2160$) were selected from this dataset, averaging 81.3 minutes. These videos capture city tours from an egocentric perspective and include 65 questions. The mean GT time interval is 31.4 seconds.
@@ 9.3 Localization evaluation
FALCON-Bench requires models to provide an evidence of the answer (short clip) rather than precisely matching the entire clip where the answer is located. To achieve this goal, we leverage The Ground Truth over Union (GToU) metric to compare the predicted and the GT temporal interval. Unlike the commonly used Intersection over Union (IoU) in temporal grounding tasks [16], GToU assigns a score of 1.0 if the predicted interval is entirely within the GT interval, regardless of the degree of overlap. In all other cases, GToU behaves identically to IoU (Figure 5). In mathematical terms,
@@ 11.1 FALCON-Bench
mPLUG-Owl3 [58] 57.6 39.2 21.1 26.0 28.7 2.24 1.20 1.58 1.67
LLaVA-OV [22] 72.2 45.0 28.5 46.0 39.8 2.46 1.57 2.42 2.15
@@
VideoChat2-HD [25] 27.2 18.6 12.1 22.0 17.5 1.24 0.67 1.34 1.08
Video-LLAVA [26] 34.0 19.6 6.49 18.0 14.6 1.20 0.46 1.14 0.93
LLaVA-Video [23] 79.2 61.7 44.0 52.0 52.6 3.31 2.38 2.74 2.81
@@
MovieChat-OV [45] 44.5 5.88 18.3 14.0 12.7 0.49 1.02 0.82 0.77
Apollo [70] 71.3 50.9 31.3 48.0 43.4 2.77 1.79 2.62 2.39
@@
mPLUG-Owl3 [58] 82.9 59.8 53.1 58.0 65.2 3.25 2.89 3.18 3.10
LLaVA-OV [22] 84.4 40.1 10.4 54.0 34.8 2.02 0.53 2.80 1.78
@@ 11.3 VLMs Calibration
To measure VLM calibration, we adopt the Reliability Diagrams [13]. These diagrams group all the predictions in bins according to their confidence and measure the gap between confidence and accuracy per each bin. Specifically, we split the $N$ predictions in $M=10$ bins according to their confidence and, for each bin $B_{m}$, we compute its count $N_{m}$ average confidence $C_{m}$ and its average accuracy $A_{m}$. From these diagrams, we may compute the Average Calibration Error (ACE) as,
@@
where $M^{+}$ is the number of non-empty bins [34]. We compute Calibration Count (CC), which quantifies the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE). For example, the Calibration Count at threshold 0.9 is computed as,
@@
We evaluated various confidence aggregation methods, specifically likelihood, average, and geometric average, as done in similar calibration studies with LLMs [30]. Looking first at the reliability diagrams (Figure 10 (a-f)), we discard likelihood as a suitable confidence metric. The distribution of confidence values is highly spread out, with many answers assigned extremely low confidence. This occurs because likelihood multiplies the probability of all tokens, thus longer answers tend to receive lower confidence scores, making it unreliable for calibration.
@@ 11.5 FALCON-Bench vs QAEgo4D
The QAEgo4D [5] dataset is the closest to FALCON-Bench in terms of the task definition as it also addresses the problem of VAS, that the authors call question answering visual language grounding (VLG). They also have open ended questions. However, besides the time differences in the video durations –FALCON-Bench with an average duration of $\sim 80$ minutes while QAEgo4D has an average duration of $\sim 8$ minutes–, the type of questions are completely different. QAEgo4D questions are automatically generated from the sparse video narration, focusing on the main object or action: Q: What did I Put in the Pan?A: cheese, or Q: What paint can did I open?A:black paint. Meanwhile, FALCON-Bench questions are human curated to be challenging: Q: What is the number of the train that crosses paths with Lightning McQueen at night?A:A113; Q:What message about the flu appears on a city building?A:Bovril nourishes you to resist ’flu or Q:Which player of Chelsea received a red card?A: Thibaut Courtois.